I need help solving the exponential equation
e 2-x=12

Answer by stanbon(75887)

e^(2-x) = 12
Take the natural log of both sides to get:
2-x = ln(12)
x = 2-ln(12)
x = -0.48491...
Cheers,
Stan H.